OpenStudy (anonymous):

A certain drug is made from only two ingredients: compound A and compound B. There are 6 milliliters of compound A used for every 5 milliliters of compound B. If a chemist wants to make 957 milliliters of the drug, how many milliliters of compound A are needed?

OpenStudy (anonymous):

I believe you are missing the number values for the amount of milliliters for each compound

OpenStudy (anonymous):

fixed it

jimthompson5910 (jim_thompson5910):

A/B = 6/5 A = (6/5)B ------------------- A+B = 957 (6/5)B + B = 957 (6/5)B + (5/5)B = 957 (11/5)B = 957 11B = 957*5 11B = 4785 B = 4785/11 B = 435 Use this to find the value of A